What capacitor in series with a 100 Ohm resistor a
Answer #1

1. Resonance frequency for series RLC circuit

  f = \frac{1}{2\pi \sqrt{LC}}

C = \frac{1}{4\pi ^{2}f^{2}L}=\frac{1}{4\pi ^{2}(800)^{2}(20\times 10^{-3})}=1.2F

2.

Inductor impedence = 2\pi fL=2\pi (60)(0.130)=49.0

capacitor impedence = -j/2\pi fC=-j/(2\pi (60)(20.0\times 10^{-6}))=132.6

Total impedence Z = R+2\pi fL-j/2\pi fC= 99.0+49.0-132.6=15.4\Omega

Peak current

I = V/Z= 120V/15.4\Omega =7.79A

3.

C = \frac{1}{4\pi ^{2}f^{2}L}=\frac{1}{4\pi ^{2}(105.6)^{2}(0.2\times 10^{-6})}=11.35F

4.

R1 becomes brighter. Connecting a wire from a to b provides a nearly zero resistance path from b to c and decreases the total resistance of the circuit from R1 + R2 to just R1. Ignoring internal resistance, the potential difference maintained by the battery is unchanged while the resistance of the circuit has decreased. The current passing through bulb R1 increases, causing this bulb to glow brighter. Bulb R2 goes out because essentially all of the current now passes through the wire connecting a and b and bypasses the filament of Bulb R2.

5.

Average power dissipated in the resistor circuit during one cycle

P =\frac{E_{rms}^{2}}{R}

It does not change with the frequency

6.

\varepsilon =-L\frac{\Delta I}{\Delta t}=-(N\phi _{B}/I)\left ( \frac{\Delta I}{\Delta t} \right )

\phi =-\frac{0.5(10)}{(1500)(130)}=-2.56\times 10^{-5}Wb